OpenStudy (anonymous):

On the graph of the equation 2x + 5y = –30, what is the value of the y-intercept?

OpenStudy (anonymous):

put x=0 and solve for y

OpenStudy (anonymous):

2(0)+5y=-30 y=-30/5 y=-6 then what ?

OpenStudy (anonymous):

yes. thats it. -6 is the value of y intercept of the given equation.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

welcome